OpenStudy (anonymous):

Winter heating of an office building requires an average of 270 kBtu/h. The heating system uses a heat pump with COP = 3.0. If the energy supplied to the heat pump costs $0.07 kWh, what's the cost per day to heat the building?

OpenStudy (anonymous):

Read this for details http://en.wikipedia.org/wiki/Coefficient_of_performance from COP=Q/W 3=270000/W W = 90000 = 90 kW And because day on planet earth is 24 hours 0.07*24*90= your cost per day = 151.2 $ (I might be wrong at this)